### v1

@@ Abstract
In knowledge-intensive tasks, especially in high-stakes domains like medicine and law, it is critical not only to retrieve relevant information but also to provide causal reasoning and explainability. Large language models (LLMs) have achieved remarkable performance in natural language understanding and generation tasks. However, they often suffer from limitations such as difficulty in incorporating new knowledge, generating hallucinations, and explaining their reasoning process. To address these challenges, integrating knowledge graphs with Graph Retrieval-Augmented Generation (Graph RAG) has emerged as an effective solution. Traditional Graph RAG methods often rely on simple graph traversal or semantic similarity, which do not capture causal relationships or align well with the model’s internal reasoning steps. This paper proposes a novel pipeline that filters large knowledge graphs to emphasize cause-effect edges, aligns the retrieval process with the model’s chain-of-thought (CoT), and enhances reasoning through multi-stage path improvements. Experiments on medical question-answering tasks show consistent gains, with up to a 10% absolute improvement across multiple large language models (LLMs). This approach demonstrates the value of combining causal reasoning with stepwise retrieval, leading to more interpretable and logically grounded solutions for complex queries.
@@ 1 Introduction
Human civilization has always advanced by systematically gathering and applying knowledge. In contemporary settings, this process typically involves integrating diverse information from large text corpora, databases, and real-world observations. Domains such as healthcare, law, and scientific research illustrate the increasing complexity of this task: clinicians correlate patient records, guidelines, and published findings to make diagnostic decisions; legal professionals cross-reference statutes, precedents, and factual evidence to build robust arguments; and scientists synthesize large volumes of experimental data to propose novel hypotheses. The ability to retrieve, analyze, and reason across multiple sources—collectively described as knowledge-intensive queries—is essential for achieving precision, clarity, and innovation [1, 2].
1.1 From Semantic Association to Causal Reasoning: Retrieval Challenges in Knowledge-Intensive Queries
@@
Significant progress has been made using retrieval-augmented paradigms where queries are matched to semantically related documents [3, 4]. However, many knowledge-intensive tasks demand more than semantic overlap: they require understanding logical structure and causal pathways that connect pieces of evidence [5]. For instance, retrieving an article stating “Disease X is associated with Gene Y” may be thematically relevant but might not fully answer how or why Gene Y leads to a specific disease manifestation. Hence, purely correlation-based retrieval can fall short in scenarios where an explicit explanation or directional link is needed.
@@
Moreover, large knowledge graphs (KGs) can exacerbate retrieval complexity. As the number of entities and edges grows—often mixing correlational or loosely defined relationships—simple similarity-based traversals may retrieve extensive sets of partially relevant or conflicting facts [6, 7]. Systems relying on dense or symbolic retrieval alone can thus amass large volumes of loosely associated data, lacking a robust mechanism for highlighting causal edges that yield a coherent inference path, frequently resulting in incomplete or contradictory answers.
@@
Recent large language models (LLMs), exemplified by GPT-4 [8] and Claude 3 [9], have demonstrated remarkable capabilities in natural language generation and single-step knowledge retrieval. Yet, these models often struggle to maintain consistent reasoning across multiple inference steps, which can lead to hallucinations or logically inconsistent conclusions [11]. Without a transparent mechanism to explain why certain pieces of information connect to others, LLM-based pipelines cannot reliably ensure that each retrieval step aligns with an evolving chain of reasoning [12]. This situation makes it difficult to confirm the correctness of multi-stage answers and raises interpretability and reliability concerns.
@@
An emerging consensus holds that causality is a critical missing piece in knowledge-intensive tasks [13, 14]. Correlation-based retrieval (e.g., “A relates to B”) may suffice for thematically aligned lookups but fails to encode directional or explanatory structures. By contrast, causal relationships indicate how and why one factor influences another [15], an especially crucial requirement in medical, legal, and scientific problem-solving.
@@
1.2 Key Limitations in Existing Graph RAG Approaches
While graph-based retrieval-augmented generation (RAG) has delivered improvements in various multi-hop tasks, the following interlocking challenges continue to undermine its reliability and explain ability—especially in complex knowledge domains such as healthcare:
@@ 1.2 Key Limitations in Existing Graph RAG Approaches
    Lack of Explicit Causal Priorities: Many existing systems treat all edges as equally important or relevant, overshadowing the truly causal links amid a sea of purely semantic or correlational relations [7]. These relations lack directionality or explanatory power.
    Effect: Because the retrieval engine cannot distinguish strong cause-effect edges from weaker associative edges, the final retrieval often includes tangential or trivial facts. As a result:
    • Path Confusion: The system’s final multi-hop path may jump between loosely connected nodes, making it hard for the model to isolate the true mechanism underlying the question.
@@
    • Reduced Explain Ability: Without focusing on cause-effect relationships, the model’s explanations (or chain-of-thought) become generic or incoherent, hampering trust in the system for high-stakes domains like healthcare.
@@
    Effect: This misalignment severely undermines the synergy between how the model thinks and which facts are fetched:
@@
    Effect: The sheer scale and connectivity of the KG can lead to:
    • Excessive Graph Noise: With so many edges often correlational or tangential—the system must filter out large quantities of low-utility data. If the filtering is inadequate or too lenient, the correct cause-effect path is buried under irrelevant expansions.
@@
Addressing these aspects is crucial for delivering reliable multi-hop QA especially in domains like medicine, where even small misalignments or spurious edges can produce dangerously incorrect recommendations. In the next sections, the proposed Causal Graph approach is presented, combined with chain-of-thought–driven retrieval, to tackle these core limitations for more trustworthy and explainable knowledge-intensive reasoning.
1.3 Hierarchical Causal Retrieval and Chain-of-Thought Enhancement
In this framework, two principles hold equal importance: (1) prioritizing causal edges within a large knowledge graph, and (2) aligning retrieval with the LLM’s chain-of-thought to dynamically guide each reasoning step. Concretely, a hierarchical retrieval paradigm is proposed that looks for strong cause-effect links first and resorts to broader correlational or semantic edges only if purely causal paths do not suffice. Termed Causal-First Graph RAG, this approach ensures that potential cause-effect chains are explored early, while still preserving a fallback route for coverage.
  1. 1.
    Minimizes extraneous retrieval: By emphasizing the graph’s most causally relevant links first, the pipeline reduces the volume of loosely connected or irrelevant nodes and edges commonly returned by correlation-based queries.
  2. 2.
    Retains broader edges if needed: If no high-confidence causal path adequately addresses a query, the pipeline can incorporate less certain or associative edges for coverage.
  3. 3.
    Chain-of-thought synergy: Through step-by-step prompting, the model’s intermediate reasoning states (“->” segments) explicitly steer retrieval, revealing which links or nodes are most essential at each inference hop.
By melding a causal-first search strategy with chain-of-thought alignment, the system targets logically grounded answers, avoiding the pitfalls of unbounded or purely correlation-based graph traversals.
@@ 1.3 Hierarchical Causal Retrieval and Chain-of-Thought Enhancement
This work presents a Causal and Chain-of-Thought Enhancement on Graph RAG framework that:
@@
    Applies a causal inference–driven filtering stage to remove contradictory or tangential sub-paths, retaining only the subgraph consistent with the model’s chain-of-thought.
Taken together, these contributions advance both retrieval quality (through causal prioritization) and reasoning consistency (through CoT guidance), culminating in a scalable pipeline for complex, multi-hop QA in medical and similarly knowledge-intensive domains.
@@ 2.1 Retrieval-Augmented Generation and Graph-Based Knowledge Discovery
Retrieval-Augmented Generation (RAG) [19] has emerged as a central paradigm for embedding external knowledge into generative models, thereby enhancing their performance on knowledge-intensive tasks. Traditional text-based retrieval pipelines—often realized via dense passage retrieval (DPR) [20, 21]—locate passages closely matching the query’s semantics, then provide them to the language model for answer generation or content synthesis. Although this approach offers clear advantages in single-fact lookups or short-answer scenarios, it may falter when queries involve complex knowledge integration or require coherent logical structure [22, 4].
To more systematically encode relational information, a number of studies advocate graph-based retrieval methods, wherein knowledge graphs (KGs) store explicit connections among entities [23, 6]. Graph RAG systems traverse these links to assemble more directed and interpretable evidence sets. This strategy supports entity disambiguation, relational reasoning, and targeted knowledge selection, providing a structured alternative to raw text passages. However, most present-day graph retrieval systems rely heavily on correlation-based edges—such as “associated with” or “related to”—that do not necessarily capture the directionality or causal significance required by more demanding queries [7, 18]. As a result, systems can retrieve relevant but non-explanatory connections, yielding incomplete or logically inconsistent knowledge sets in scenarios that demand deeper causal or sequential reasoning.
@@ 2.2 Causality in Knowledge-Intensive Natural Language Processing
Causality has long been recognized as a fundamental cornerstone of advanced knowledge discovery in fields spanning cognitive science, epidemiology, and machine learning [17, 24]. Within natural language processing (NLP), numerous lines of inquiry—from causal event extraction [25] and counterfactual generation to specialized question answering—aim to move beyond simple correlations toward why-oriented explanations [13, 14]. Nevertheless, integrating explicit causal inference into retrieval-centric architectures remains nontrivial. Many frameworks stop at correlation-based links or unstructured text retrieval, lacking specialized graph structures or annotated metadata to differentiate cause-and-effect from adjacency or co-occurrence [15]. Hence, while semantically relevant facts can be extracted, the logical consistency or directional flow crucial in knowledge-intensive queries is not guaranteed [12].
@@ 2.3 Chain-of-Thought Prompts and Reasoning Transparency
Chain-of-thought (CoT) prompting [16] has recently garnered attention as a method to expose a language model’s intermediate reasoning path. By prompting the model to sequentially articulate partial conclusions rather than jumping to a final answer, CoT can (i) provide interpretable glimpses into the model’s decision-making process and (ii) sometimes improve performance on tasks requiring compositional or multi-stage thinking [1]. However, the widespread usage of CoT remains largely decoupled from retrieval pipelines. Typically, the LLM generates an internal chain-of-thought, but the system does not subdivide retrieval according to each CoT step. Consequently, any fine-grained reasoning states are not matched with equally fine-grained retrieval steps, creating potential mismatches.
@@ 2.4 Challenges in Aligning Causality, Graph Retrieval, and CoT Reasoning
    Causal Under-specification. Many graph-based retrieval modules store correlations or broad associations but do not designate strong cause-and-effect edges. As a result, the pipeline fails to differentiate essential causal links from incidental relationships [7].
@@
    Misaligned Retrieval vs. Reasoning Steps. Chain-of-thought prompts typically divide the reasoning into multiple short sub-steps, but standard Graph RAG retrieval is organized as a single module that either returns a large chunk of data or lumps all evidence together [12]. This misalignment can cause the LLM to request step-by-step clarifications, yet the retrieval engine has not subdivided the knowledge base lookups accordingly, thus returning content irrelevant to the current sub-question.
@@
    Scalability Issues with Large Knowledge Graphs. In real-world applications, knowledge graphs may contain millions of edges with varying specificity or utility [6]. A naive graph traversal often yields excessively large or tangentially related subgraphs, eclipsing the critical edges. This surplus can confuse both causal reasoning and chain-of-thought, resulting in contradictory or incomplete final answers.
The present study introduces a unified framework that elevates causal edges within a large-scale graph environment while synchronizing retrieval with the LLM’s chain-of-thought states. By assigning higher priority to cause-effect relationships in the subgraph search, the pipeline avoids overloading the model with loosely associated edges. Furthermore, each chain-of-thought step prompts a targeted query, ensuring fine-grained alignment between what the LLM considers at a given moment and which graph edges or nodes are returned. This synergy preserves the interpretability gains of CoT while injecting causal rigor into the retrieval process, thereby addressing correlation-based noise and subgraph sprawl. The subsequent sections illustrate how this hierarchical and cause-first retrieval approach can enhance knowledge-intensive QA, mitigate misalignment problems, and handle expansive KGs in a controlled, step-by-step manner.
@@ 3 Methodology
This section describes the pipeline for (1) constructing a causal subgraph from a large knowledge graph (KG), (2) generating a chain-of-thought (CoT) that segments reasoning steps using $\rightarrow$” markers, and (3) performing multi-stage retrieval and path enhancement with fallback strategies, including a final phase that re-injects the merged paths and original CoT into the language model. Figure 1 provides an overview.
Figure 1: Overall framework in three stages: (1) building/updating the causal graph, (2) chain-of-thought–driven retrieval, (3) multi-stage path processing.
@@ 3.1 Overall Framework
    Causal graph construction and updating: A large KG is scanned, and only edges carrying cause-effect significance (e.g., CAUSE, MANIFESTATION_OF) are retained, assigning numeric strengths. Edges deemed purely correlational (e.g., ASSOCIATED_WITH, RELATED_TO) are excluded if their relevance falls below a threshold.
@@
    Chain-of-thought–driven causal path retrieval: A CoT is acquired from the LLM, splitting each segment by $\rightarrow$,” and attempting to find paths in the causal subgraph that connect the relevant entities. If no match is found, the method reverts to the full KG.
@@
    Causal path processing and multi-stage enhancement: After retrieval, structurally similar paths are merged, refined scores (based on entity/semantic overlaps and path length) are computed, and the top-ranked subset is presented as final evidence. Both the final path set and the original CoT are then fed back to the LLM for a consistency check and “enhanced” output.

### v2

@@ Abstract
In knowledge-intensive reasoning tasks, especially in high-stakes domains like medicine and law, there are high demands for both the accuracy of information retrieval and the causal reasoning and explainability of the reasoning process. Large language models, due to their powerful reasoning abilities, have demonstrated remarkable potential in knowledge-intensive tasks. However, they still face challenges such as difficulty in incorporating new knowledge, generating hallucinations, and explaining their reasoning process. To address these challenges, integrating knowledge graphs with Graph Retrieval-Augmented Generation (Graph-RAG) has emerged as an effective solution. Traditional Graph-RAG methods often rely on simple graph traversal or semantic similarity, which do not capture causal relationships or align well with the model’s internal reasoning steps. This paper proposes a novel pipeline that filters large knowledge graphs to emphasize cause-effect edges, aligns the retrieval process with the model’s chain-of-thought (CoT), and enhances reasoning through multi-stage path improvements. Experiments on medical question-answering tasks demonstrate consistent improvements in diagnostic accuracy, with up to a 10% absolute gain across multiple large language models (LLMs), evaluated on two medical domain datasets. This approach demonstrates the value of combining causal reasoning with stepwise retrieval, leading to more interpretable and logically grounded solutions for complex queries.
@@ 1 Introduction
Human civilization has always advanced by systematically gathering and applying knowledge. In contemporary settings, this process typically involves integrating diverse information from large text corpora, databases, and real-world observations. Domains such as healthcare, law, and scientific research illustrate the increasing complexity of this task: clinicians correlate patient records, guidelines, and published findings to make diagnostic decisions; legal professionals cross-reference statutes, precedents, and factual evidence to build robust arguments; and scientists synthesize large volumes of experimental data to propose novel hypotheses. The ability to retrieve, analyze, and reason across multiple sources collectively known as knowledge-intensive queries is essential for achieving precision, clarity, and innovation.
1.1 Retrieval Challenges in Knowledge-Intensive Queries
@@
In Knowledge-Intensive Queries tasks, significant progress has been made using retrieval-augmented paradigms where queries are matched to semantically related documents [8, 7]. However, many knowledge-intensive tasks demand more than semantic overlap: they require understanding logical structure and causal pathways that connect pieces of evidence [17]. For instance, retrieving an article stating “Disease X is associated with Gene Y” may be thematically relevant but might not fully answer how or why Gene Y leads to a specific disease manifestation. Hence, purely correlation-based retrieval can fall short in scenarios where an explicit explanation or directional link is needed.
@@
Knowledge graphs, as an effective structural knowledge carrier, are often used in retrieval tasks. However, large knowledge graphs (KGs) can exacerbate retrieval complexity. As the number of entities and edges grows, often mixing correlational or loosely defined relationships-simple similarity-based traversals may retrieve extensive sets of partially relevant or conflicting facts [2]. Systems relying solely on dense or symbolic retrieval can thus accumulate large volumes of loosely associated data, lacking a robust mechanism for highlighting causal edges that yield a coherent inference path, frequently resulting in incomplete or contradictory answers.
@@
Recent large language models (LLMs) have demonstrated remarkable capabilities in natural language generation and single-step knowledge retrieval [1, 18]. Yet, these models often struggle to maintain consistent reasoning across multiple inference steps, which can lead to hallucinations or logically inconsistent conclusions [5]. Without a transparent mechanism to explain why certain pieces of information connect to others, LLM-based pipelines cannot reliably ensure that each retrieval step aligns with an evolving chain of reasoning [16]. This situation makes it difficult to confirm the correctness of multi-stage answers and raises interpretability and reliability concerns.
@@
An emerging consensus holds that causality is a critical missing piece in knowledge-intensive tasks [11, 13]. Correlation-based retrieval (e.g., "A relates to B") may suffice for thematically aligned lookups but fails to encode directional or explanatory structures. By contrast, causal relationships indicate how and why one factor influences another [12], an especially crucial requirement in medical, legal, and scientific problem-solving.
@@
1.2 Key Limitations in Existing Graph-RAG Approaches
While graph-based retrieval-augmented generation (RAG) has delivered improvements in various knowledge-intensive reasoning tasks [4, 3], the following interlocking challenges continue to undermine its reliability and explainability—especially in complex knowledge domains such as healthcare:
@@ 1.2 Key Limitations in Existing Graph-RAG Approaches
    Lack of Explicit Causal Priorities: Many existing systems treat all edges as equally important or relevant, overshadowing the truly causal links amid a sea of purely semantic or correlational relations. These relations lack directionality or explanatory power.
    Because the retrieval engine cannot distinguish strong cause-effect edges from weaker associative edges, the final retrieval often includes tangential or trivial facts. As a result:
    • Path Confusion: The system’s final reasoning path may jump between loosely connected nodes, making it hard for the model to isolate the true mechanism underlying the question.
@@
    • Reduced Explainability: Without focusing on cause-effect relationships, the model’s explanations (or chain-of-thought) become generic or incoherent, hampering trust in the system for high-stakes domains.
@@
    This misalignment severely undermines the synergy between how the model thinks and which facts are fetched:
@@
    The sheer scale and connectivity of the KG can lead to:
    • Excessive Graph Noise: With so many edges often correlational or tangential, the system must filter out large quantities of low-utility data. If the filtering is inadequate or too lenient, the correct cause-effect path is buried under irrelevant expansions.
@@
Addressing these aspects is crucial for delivering reliable knowledge-intensive reasoning tasks especially in domains like medicine, where even small misalignments or spurious edges can produce dangerously incorrect recommendations.
@@
In this framework, two principles hold equal importance: (1) prioritizing causal edges within a large knowledge graph, and (2) aligning retrieval with the LLM’s chain-of-thought to dynamically guide each reasoning step. Concretely, a hierarchical retrieval paradigm is proposed that looks for strong cause-effect links first and resorts to broader correlational or semantic edges only if purely causal paths do not suffice. Termed Causal-First Graph-RAG, this approach ensures that potential cause-effect chains are explored early, while still preserving a fallback route for coverage.
The framework of this research is as follows:
@@
    Applies a causal inference-driven filtering stage to remove contradictory or tangential sub-paths, retaining only the subgraph consistent with the model’s chain-of-thought.
Taken together, these contributions advance both retrieval quality (through causal prioritization) and reasoning consistency (through CoT guidance), culminating in a scalable pipeline for complex, knowledge-intensive reasoning tasks in medical and similarly knowledge-intensive domains.
@@ 2.1 Retrieval-Augmented Generation and Graph-Based Knowledge Discovery
Retrieval-Augmented Generation (RAG) [8] has become a central paradigm for embedding external knowledge into generative models, improving performance on knowledge-intensive tasks. Traditional text-based retrieval, such as dense passage retrieval (DPR) [7], is effective for single-fact lookups but struggles with queries requiring complex knowledge integration or logical coherence. To more systematically encode relational information, a number of studies advocate graph-based retrieval methods, wherein knowledge graphs (KGs) store explicit connections among entities. Graph-RAG systems traverse these links to assemble more directed and interpretable evidence sets. This strategy supports entity disambiguation, relational reasoning, and targeted knowledge selection, providing a structured alternative to raw text passages. However, most present-day graph retrieval systems rely heavily on correlation-based edges—such as "associated with" or "related to"—that do not necessarily capture the directionality or causal significance required by more demanding queries. As a result, systems can retrieve relevant but non-explanatory connections, yielding incomplete or logically inconsistent knowledge sets in scenarios that demand deeper causal or sequential reasoning.
@@ 2.2 Causality in Knowledge-Intensive Natural Language Processing
Causality has long been recognized as a fundamental cornerstone of advanced knowledge discovery in fields spanning cognitive science, epidemiology, and machine learning [10]. Within natural language processing (NLP), numerous lines of inquiry—from causal event extraction and counterfactual generation to specialized question answering—aim to move beyond simple correlations toward why-oriented explanations [11, 13]. Nevertheless, integrating explicit causal inference into retrieval-centric architectures remains nontrivial. Many frameworks stop at correlation-based links or unstructured text retrieval, lacking specialized graph structures or annotated metadata to differentiate cause-and-effect from adjacency or co-occurrence [12]. Hence, while semantically relevant facts can be extracted, the logical consistency or directional flow crucial in knowledge-intensive queries is not guaranteed.
@@ 2.3 Chain-of-Thought Prompts and Reasoning Transparency
Chain-of-thought (CoT) prompting [16] has recently garnered attention as a method to expose a language model’s intermediate reasoning path. By prompting the model to sequentially articulate partial conclusions rather than jumping to a final answer, CoT can (i) provide interpretable glimpses into the model’s decision-making process and (ii) sometimes improve performance on tasks requiring compositional or multi-stage thinking. However, the widespread usage of CoT remains largely decoupled from retrieval pipelines. Typically, the LLM generates an internal chain-of-thought, but the system does not subdivide retrieval according to each CoT step. Consequently, any fine-grained reasoning states are not matched with equally fine-grained retrieval steps, creating potential mismatches.
@@ 2.4 Challenges in Aligning Causality, Graph Retrieval, and CoT Reasoning
    Causal Under-specification. Many graph-based retrieval modules store correlations or broad associations but do not designate strong cause-and-effect edges. As a result, the pipeline fails to differentiate essential causal links from incidental relationships.
@@
    Misaligned Retrieval vs. Reasoning Steps. Chain-of-thought prompts typically divide the reasoning into multiple short sub-steps, but standard Graph-RAG retrieval is organized as a single module that either returns a large chunk of data or lumps all evidence together. This misalignment can cause the LLM to request step-by-step clarifications, yet the retrieval engine has not subdivided the knowledge base lookups accordingly, thus returning content irrelevant to the current sub-question.
@@
    Scalability Issues with Large Knowledge Graphs. In real-world applications, knowledge graphs may contain millions of edges with varying specificity or utility. A naive graph traversal often yields excessively large or tangentially related subgraphs, eclipsing the critical edges. This surplus can confuse both causal reasoning and chain-of-thought, resulting in contradictory or incomplete final answers.
The present study introduces a unified framework that elevates causal edges within a large-scale graph environment while synchronizing retrieval with the LLM’s chain-of-thought states. By assigning higher priority to cause-effect relationships in the subgraph search, the pipeline avoids overloading the model with loosely associated edges. Furthermore, each chain-of-thought step prompts a targeted query, ensuring fine-grained alignment between what the LLM considers at a given moment and which graph edges or nodes are returned. This synergy preserves the interpretability gains of CoT while injecting causal rigor into the retrieval process, thereby addressing correlation-based noise and subgraph sprawl.
@@ 3 Methodology
This section describes the pipeline for (1) constructing a causal subgraph from a large knowledge graph (KG), (2) generating a chain-of-thought (CoT) that segments reasoning steps using $\rightarrow$" markers, and (3) performing multi-stage retrieval and path enhancement with fallback strategies, including a final phase that re-injects the merged paths and original CoT into the language model. Figure 1 provides an overview.
Figure 1: Overall framework in three stages: (1) building/updating the causal graph, (2) chain-of-thought-driven retrieval, (3) multi-stage path processing.
@@ 3.1 Overall Framework
    Causal graph construction and updating: A large KG is scanned, and only edges carrying cause-effect significance are retained, assigning numeric strengths. Edges deemed purely correlational are excluded if their relevance falls below a threshold.
@@
    Chain-of-thought-driven causal path retrieval: A CoT is acquired from the LLM, splitting each segment by $\rightarrow$," and attempting to find paths in the causal subgraph that connect the relevant entities. If no match is found, the method reverts to the full KG.
@@
    Causal path processing and multi-stage enhancement: After retrieval, structurally similar paths are merged, refined scores (based on entity/semantic overlaps and path length) are computed, and the top-ranked subset is presented as final evidence. Both the final path set and the original CoT are then fed back to the LLM for a consistency check and "enhanced" output.